\begin{thebibliography}{9}\footnotesize\itemsep-2pt
\bibitem{trianglefree2}
  {\sc A. Al-Azemi, A. Betten},
  {\it Classification of triangle-free $22_3$ configurations},
  Int. J. Comb. {\bf 2010}(2010), Art. Id. 767361, 1--17.
\bibitem{betten}
  {\sc D. Betten, U. Schumacher},
  {\it The ten configurations ${10}_{3}$},
  Rostock. Math. Kolloq.
  {\bf 46} (1993), 3--10.
\bibitem{trianglefree1}
  {\sc M. Boben, B. Gr{\"u}nbaum, T. Pisanski, A. Zitnik},
  {\it Small triangle-free configurations of points and lines},
  Discrete. Comput. Geom. {\bf 35} (2006), 405--427.
\bibitem{doliwa1}
  {\sc A. Doliwa},
  {\it The affine Weil group symmetry of Desargues maps and the non-commutative
  Hirota-Miwa system},
  Phys. Lett. A {\bf 375} (2011), 1219--1224.
\bibitem{doliwa2}
  {\sc A. Doliwa},
  {\it Desargues maps and the Hirota-Miwa equation},
  Proc. R. Soc. A {\bf 466} (2010), 1177--1200.
\bibitem{6sparse}
  {\sc A. D. Forbes, M. J. Granell, T. S. Griggs},
  {\it On $6$-sparse Steiner triple systems},
  J. Combin. Theory Ser. A {\bf 144}(2007), no .2, 235--252.
\bibitem{antimitre2}
  {\sc Y. Fuiiwara},
  {\it Infinite classes of anti-mitre and $5$-sparse Steiner triple systems},
  J. Combin. Des. {\bf 14} (2006), no. 3, 237--250.
\bibitem{veradjac}
  {\sc I. Golonko, M. Pra{\.z}mowska},
  {\it Adjacency in Generalized Projective Veronese Spaces},
  Abh. Math. Sem. Univ. Hamb. {\bf 76}(2006), 99--114.
\bibitem{grop1}
  {\sc H. Gropp},
  {\it Configurations and their realization},
  {\it Combinatorics (Rome and Montesilvano, 1994)},
  Discrete Math. {\bf 174}, no. 1--3 (1997), 137--151.
\bibitem{trianglefree3}
  {\sc B. Gr{\"u}nbaum},
  {\it Geometric realization of some triangle-free combinatorial configurations
  $22_3$},
  ISRN Geom {\bf 2012} (2012), ArticleID 560760, 1--10.
\iffalse
\bibitem{grunbaum}
  {\sc B. Gr{\"u}nbaum},
  {\it Configurations of points and lines},
  American Mathematical Society, Providence, RI, {\bf 103} (2009),
\fi
%% \bibitem{hilbert}
%%   {\sc D. Hilbert, P. Cohn-Vossen},
%%   {\sl Anschauliche Geometrie},
%%   Springer Verlag, Berlin, 1932.
%%  (English translation: {\it Geometry and the Imagination}, AMS Chelsea Publishing).
\bibitem{kantor}
  {\sc S. Kantor},
  {\it Die Konfigurationen $(3,3)_{10}$},
  Sitzungsberichte Wiener Akad. {\bf 84} (1881), 1291--1314.
\bibitem{klimczak}
  {\sc A. Klimczak, M. Pra{\.z}mowska},
  {\it Desarguesian closure of binomial graphs},
  Demonstratio Math. {\bf 39}(2006), no. 2 245--253
\iffalse
\bibitem{levi}
  {\sc F. Levi},
  {\it Geometrische Konfigurationen},
  Leipzig, S. Hirzel (1929)
\fi
\bibitem{antipasch1}
  {\sc A. C. H. Ling, C. J. Colbourn, M. J. Granell, T. S. Griggs},
  {\it Construction techniques for anti-Pasch Steiner triple systems},
  Jour. London Math. Soc. {\bf 61} (2000), no. 3, 641--657.
\iffalse
\bibitem{mathon}
  {\sc R.A. Mathon, K.T. Phelps, A. Rosa},
  {\it Small Steiner triple systems and their properties},
  Ars Combin. {\bf 15} (1983), 3--110; 
Errata: 
{\it Small Steiner triple systems and their properties},
  Ars Combin. {\bf 16} (1983), 286.
\fi
%% \bibitem{petel}
%%  {\sc K. Petelczyc},
%%  {\it Series of inscribed $n$-gons and rank 3 configurations},
%%  Contributions to Alg. and Geom. {\bf 46} (2005), No.\;1, 283-300  
\bibitem{linstp}
  {\sc K. Petelczyc, M. Pra{\.z}mowska},
  {\it Twisted Fano spaces and their classification, linear completions of systems of % 
  triangle perspectives},
  Des. Codes Cryptogr. {\bf 54} (2010), 241--251. 
\bibitem{klik:VC}
  {\sc K. Petelczyc, M. Pra{\.z}mowska},
  {\it ${10}_{3}$-configurations and projective realizability of multiplied configurations},
  Des. Codes Cryptogr. {\bf 51}, no. 1 (2009), 45--54. 
%% \bibitem{pikert}
%%   {\sc G. Pickert},
%%   {\it Projektive Ebenen},
%%   Springer Verlag, Berlin 1975.
\bibitem{kwardfree}
  {\sc T. Pisanski, M. Boben, D. Maru$\hat{\text{s}}$i$\hat{\text{c}}$, A. Orbani{\'c},
  A. Graovac},
  {The $10$-cages and derived configurations},
  Discrete Math. {\bf 257} (2004), 265--276.
\iffalse
\bibitem{obrazki}
  {\sc B. Polster},
  {\sl A geometrical picture book},
  Springer Verlag, New York, 1998.
\fi
%% \bibitem{twistfan}
%%   {\sc M. Pra{\.z}mowska},
%%   {\it Twisted projective spaces and linear completions of some %
%%   partial Steiner triple systems},
%% %%  Communications Algebra Geom., 
%%   Beitr{\"a}ge Algebra Geom.,
%%   {\bf 49} no. 2 (2008), 341--368.
\bibitem{combver}
  {\sc M. Pra{\.z}mowska, K. Pra{\.z}mowski},
  {\it Combinatorial Veronese structures, their geometry, and problems of %
  embeddability},
  Results Math. {\bf 51} (2008), 275--308.
\bibitem{pascvebl}
  {\sc M. Pra{\.z}mowska, K. Pra{\.z}mowski},
  {\it Some generalization of Desargues and Veronese configurations},
  Serdica Math. J. {\bf 32} (2006), no {2--3}, {185--208}.
\bibitem{perspect}
  {\sc M. Pra{\.z}mowska},
  {\it Multiple perspectives and generalizations of the Desargues configuration},
  Demonstratio Math. {\bf 39} (2006), no. 4, {887--906}.
\bibitem{mveb2proj}
  {\sc M. Pra{\.z}mowska},
  {\it On the existence of projective embeddings of some multiveblen configurations},
  Bull. Belg. Math. Soc. Simon-Stevin {\bf 17} (2010), 1--15.
\bibitem{skewgras}
  {\sc M. Pra{\.z}mowska},
  {\it On some regular multi-veblen configurations, the geometry of 
  combinatorial quasi Grassmannians},
  Demonstratio Math. {\bf 42} (2009), no.2 387--402.
\bibitem{steiniz}
  {\sc E. Steinitz},
  {\it Konfigurationen der projektiven Geometrie},
  Encyclop. d. math. Wiss.  III (Geometrie) 1910, 481--516.
\bibitem{antimitre1}
  {\sc A. Wolfe},
  {\it The resolution of the antimitre Steiner triple system conjecture},
  J. Combin. Des. {\bf 14}(2006), no. 3, 229--236.
\end{thebibliography}

\documentclass[a4paper, 11pt, underruledhead]{paper}
\usepackage{math}
\usepackage[mathscr]{euscript}
\usepackage{enumerate}
\newif\ifnoweb
\nowebtrue

\begin{document}



\let\goth\mathfrak

\def\myend{{}\hfill{\small$\bigcirc$}}

\newenvironment{ctext}{%
  \par
  \smallskip
  \centering
}{%
 \par
 \smallskip
 \csname @endpetrue\endcsname
}

\def\id{\mathrm{id}}

\newcommand{\msub}{\mbox{\large$\goth y$}}    
\def\suport{{\mathrm{supp}}}
\newcommand{\nat}{{N}}   
\def\bagnom#1#2{\genfrac{[}{]}{0pt}{}{#1}{#2}}

\def\kros(#1,#2){c_{\{ #1,#2 \}}}
\def\skros(#1,#2){{\{ #1,#2 \}}}
\def\LineOn(#1,#2){\overline{{#1},{#2}\rule{0em}{1,5ex}}}
\def\inc{\mathrel{\strut\rule{3pt}{0pt}\rule{1pt}{9pt}\rule{3pt}{0pt}\strut}}
\def\lines{{\cal L}}
\def\tran{{\mathscr T}}
\def\collin{\sim}
\def\wspolin{{\bf L}}
\def\ncollin{\not\collin}

\def\VerSpace(#1,#2){{\bf V}_{{#2}}({#1})}
\def\GrasSpace(#1,#2){{\bf G}_{{#2}}({#1})}
\def\VeblSpSymb{{\bf P}\mkern-10mu{\bf B}}
\def\VeblSpace(#1){\VeblSpSymb({#1})}
\def\xwpis#1#2{{{\triangleright}\mkern-4mu{{\strut}_{{#1}}^{{#2}}}}}
\def\MultVeblSymb{{\sf M}\mkern-10mu{\sf V}}        
\def\MultVeblSymq{{\sf T}\mkern-6mu{\sf P}}        
\def\xwlep(#1,#2,#3,#4){{\MultVeblSymb^{#1}{\xwpis{#2}{#3}}{#4}}}
\def\xwlepp(#1,#2,#3,#4,#5){{\MultVeblSymb_{#1}^{#2}{\xwpis{#3}{#4}}{#5}}}
\def\xwlepq(#1,#2,#3,#4,#5){{\MultVeblSymq_{#1}^{#2}{\xwpis{#3}{#4}}{#5}}}
\def\konftyp(#1,#2,#3,#4){\left( {#1}_{#2}\, {#3}_{#4} \right)}
\def\binokonf(#1){\konftyp({\binom{{#1}}{2}},{{#1}-2},\binom{{#1}}{3},3)}
\let\binoconf\binokonf


\newcount\liczbaa
\newcount\liczbab

\def\binkonfo(#1,#2){\liczbaa=#2 \liczbab=#2 \advance\liczbab by -2
\def\doa{\ifnum\liczbaa = 0\relax \else
\ifnum\liczbaa < 0 \the\liczbaa \else +\the\liczbaa\fi\fi}
\def\dob{\ifnum\liczbab = 0\relax \else
\ifnum\liczbab < 0 \the\liczbab \else +\the\liczbab\fi\fi}
\konftyp(\binom{#1\doa}{2},#1\dob,\binom{#1\doa}{3},3) }

\newcount\liczbac
\def\binkonf(#1,#2){\liczbac=#2 
\def\docc{\ifnum\liczbac = 0 \relax \else\ifnum\liczbac < 0\the\liczbac 
\else+\the\liczbac\fi\fi}
B_{#1{\docc}}}


\def\Ver{{\sf\bfseries VER}}
\def\Des{{\sf\bfseries DES}}
\def\Desv{{\sf\bfseries DES'}}
\def\Desr{{\sf\bfseries DES''}}
\def\MTC{{\sf STP}}
\def\MVC{{\sf MVC}}

\def\PSTS{{\sf PSTS}}
\def\BSTS{{\sf BSTS}}


\title[Binomial \PSTS's containing complete graphs]{%
Binomial partial Steiner triple systems containing complete graphs}

\author{
Ma{\l}gorzata Pra{\.z}mowska%
\footnote{{\em Corresponding Author}, e-mail: {\ttfamily malgpraz@math.uwb.edu.pl},
telephone: +48 606 74 64 56}, 
Krzysztof Pra{\.z}mowski}

\maketitle

\begin{abstract}
  We propose a new approach to studies on partial Steiner triple systems consisting
  in determining complete graphs contained in them.
  We establish the structure which complete graphs yield in a minimal \PSTS\ that
  contains them. As a by-product we introduce the notion of a binomial \PSTS\
  as a configuration with parameters of a minimal \PSTS\ with a complete subgraph.
  A representation of binomial \PSTS\ with at least a given number of its maximal
  complete subgraphs is given in terms of systems of perspectives.
  \ifnoweb
  Finally, we prove that for each admissible integer there is a binomial \PSTS\ with
  this number of maximal complete subgraphs. 
  \fi
  \begin{flushleft}
    Mathematics Subject Classification (2010): 05B30 (05C51, 05B40).\\
    Key words: binomial configuration, generalized Desargues configuration, complete graph.
  \end{flushleft}
\end{abstract}

\section*{Introduction}

In the paper we investigate the structure which (may) yield complete graphs 
contained in a (partial) Steiner triple system (in short: in a \PSTS).
Our problem is, in fact, a particular instance of a general question,
investigated in the literature, which {\sf STS}'s (more generally: which
{\sf PSTS}'s) contain/do not contain a configuration of a prescribed
type.
Here belongs, e.g. the problem to determine Pasch-free configurations
in some definite classes of ({\sf P}){\sf STS}'s  (\cite{antipasch1}), 
or to characterize {\sf STS}'s
which do not contain the mitre configuration (e.g.: \cite{antimitre1}
\cite{antimitre2})). 
It resembles also (and, in a sense, generalizes) the
problem to 
determine triangle-free (e.g. \cite{trianglefree2}, \cite{trianglefree3}), 
quadrangle-free (e.g. \cite{trianglefree1}), 
and other $\cal K$-free (e.g. \cite{6sparse}, \cite{kwardfree})
configurations, where $\cal K$ is a fixed class of configurations (of a
special importance).
In our case configurations in question are 
complete graphs and so-called binomial configurations.

In essence, in what follows, speaking about a (complete) graph $G$ contained in
a configuration $\goth M$ we always assume that $G$ is {\em freely} contained 
in $\goth M$ i.e. it is not merely a subgraph of 
the collinearity (point-adjacency) graph of $\goth M$ 
but also distinct edges of $G$ lie on distinct lines (so called {\em sides})
of $\goth M$, and sides do not intersect outside $G$.
Note that following this terminology a complete quadrangle on a projective plane $\goth P$
{\em is not} a $K_4$-graph {\em contained} in $\goth P$.

\medskip
There are two main problems that this theory starts with.
Firstly: 
{\em what are the minimal parameters of a \PSTS\ necessary to contain a $K_n$-graph}
and 
{\em does there exist  a \PSTS\ with these parameters which contains/does not contain
a $K_n$-graph}.
It turns out that a minimal (with respect to its size) \PSTS\ that contains $K_n$ is a
so called {\em binomial configuration} i.e. a $\konftyp(v,r,b,3)$-configuration
such that
$v =\binom{m}{2}$, $r = m-2$, and $b = \binom{m}{3}$ for a positive integer $m$ ($=n+1$)
(Prop. \ref{minSTS2pelny}).
We say, in short, that 
this \PSTS\ is {\em a minimal configuration} which contains $K_n$.

Several classes of binomial configurations were introduced and studied in the literature
(see \cite[combinatorial Grassmannians]{perspect}, 
\cite[combinatorial Veronesians]{combver},
\cite[multiveblen configurations]{pascvebl}, \cite{mveb2proj}, \cite{skewgras}).
The above characterization of the parameters of a binomial configuration gives 
a good motivation and justification
for investigating this class from `a general perspective'.
In most of the binomial configurations already defined in the literature a suitable
complete graph can be found.
As the best known example of a binomial configuration
we can quote generalized Desargues configuration
(see \cite{klimczak}, \cite{doliwa1}, \cite{doliwa2}).
Then the answer to the `dual' question 
{\em what is the maximal size of a complete graph that a binomial
$\binkonfo(n,0)$-configuration $\goth M$ may contain}
easily follows: this size is $n-1$.
In this case we say that $K_n$ is a maximal (complete) subgraph of $\goth M$.
However, there are $\binkonfo(n,0)$-configurations that do not contain any $K_{n-1}$-graph.
The simplest example can be found for $n=5$: it is known that there are 
$10_3$-configurations without $K_4$.
(cf. \cite{betten}, \cite{klik:VC}).

\medskip
The second problem is read as follows: 
provided a $\binkonfo(n,1)$-configuration contains a $K_n$,
{\em what is the possible number of $K_n$-graphs contained in $\goth M$ and what is the
structure they yield}.
It turns  out that the maximal number of such $K_n$-subgraphs is $n+1$
(Prop. \ref{prop:maxK}).
Binomial configurations with the maximal number of $K_n$-subgraphs are exactly
the generalized Desargues configurations.
This fact points out once more a special position of the class of generalized 
Desargues configurations within the class of binomial configurations.

As we already mentioned,
the problem whether a given binomial $\binkonfo(n,1)$-configuration $\goth M$
contains a complete $K_n$-graph is slightly similar to the problem if a \PSTS\ contains 
a Pasch configuration. Indeed, $\goth M$ contains such a graph iff it contains
a binomial $\binkonfo(n,0)$-configuration 
(Prop. \ref{pelny2horyzont}, Cor.\ref{cor:pelny2horyzont}). 
So, (binomial) configurations 
(of size $\binom{n+1}{2}$)
without $K_n$-graphs are exactly the configurations whose no subspace is 
a binomial configuration of one-step-smaller size $\binom{n}{2}$.
\par
A binomial $\binkonfo(n,1)$-configuration with two $K_n$-subgraphs can be considered
as an abstract scheme of a perspective of two $n$-simplices
(Prop. \ref{prop:cross-compl3}).
Generally, the structure of the intersection points 
of $K_n$-subgraphs of a $\binkonfo(n,1)$-configuration
is isomorphic to a generalized Desargues configuration (Prop. \ref{prop:maxK}).

In the paper we do not intend to give a {\em detailed} analysis of the internal structure
of binomial configurations which 
contain a {\em prescribed} number of their maximal $K_n$-subgraphs.
We give only a technique (a procedure) to construct a binomial configuration which contains
at least the given number of $K_n$-subgraphs: Thm. \ref{thm:SPS}
And we prove that for each admissible integer $m$ there does exist a binomial configuration
which freely contains $m$ $K_n$-subgraphs.
Some remarks are also
made which show that, surprisingly, binomial configurations with `many' 
maximal complete subgraphs 
(the maximal admissible number, the maximal number reduced by 2, 
and the maximal number reduced by 3)
are the configurations of some well known classes.


\section{Definitions}


Recall that the term {\em combinatorial configuration} (or simply a {\em configuration}) is 
usually (cf. e.g. \cite{grop1}, \cite{steiniz}) 
used for a finite incidence point-line structure provided 
that two different points are incident with at most one line and the line size and 
point rank are constant.
Speaking more precisely, a {\em  $\konftyp(v,r,b,\varkappa)$-configuration} 
is a combinatorial configuration with 
$v$ points and $b$ lines such that
there are $r$ lines through each point, and there are $\varkappa$ points on each line.
A partial Steiner triple system (a \PSTS, in short) is a configuration 
whose each line contains exactly three points.

In what follows we shall pay special attention to so called 
{\em binomial configurations} (more precisely: binomial partial Steiner triple systems,
\BSTS, in short) i.e. to $\binkonfo(n,0)$-configurations with arbitrary integer $n\geq 2$.
The class of the configurations with these parameters will be denoted by $\binkonf(n,0)$.


Let $k$ be a positive integer and $X$ a set; we write $\sub_k(X)$ for the 
set of all $k$-subsets of $X$.
The incidence structure
\begin{ctext}
  $\GrasSpace(X,k) := \struct{\sub_k(X),\sub_{k+1}(X),\subset}$
\end{ctext}
will be called {\em a combinatorial Grassmannian} (cf. \cite{perspect}).
If $|X| = n$ then $\GrasSpace(X,2)$ is a $\binkonf(n,0)$-configuration, so it is 
a \BSTS. As the structure $\GrasSpace(X,2)$ is (up to an isomorphism) uniquely determined
by the cardinality $|X|$ in what follows we frequently write $\GrasSpace(|X|,2)$ instead
of $\GrasSpace(X,2)$. 
Recall that $\GrasSpace(4,2)$ is the Veblen (the Pasch) configuration and $\GrasSpace(5,2)$
is the Desargues Configuration (see e.g. \cite{combver}, \cite{perspect}).
Generally, every structure $\GrasSpace(n,2)$, $n\geq 5$ will be called
a {\em generalized Desargues configuration} 
(cf. \cite{skewgras}, \cite{doliwa1}, \cite{doliwa2}).

\medskip
Let $X$ be a non void set, $n = |X|$. A nondirected loopless graph
defined on $X$ (we say simply: a {\em graph})  is a structure of the form 
$\struct{X,{\cal P}}$ with ${\cal P}\subset \sub_2(X)$.
We write $K_X = \struct{X,\sub_2(X)}$ 
for the complete graph with the vertices $X$;
the term $K_n$ is used for an arbitrary graph $K_X$ where $|X| = n$.

We say that a configuration ${\goth N} = \struct{X,{\cal G}}$ 
{\em is contained in} a configuration ${\goth M} = \struct{S,\lines}$ if 
\begin{itemize}\itemsep-2pt\def\labelitemi{-}
\item
  $X \subset S$.
\item
  each line $L\in{\cal G}$ extends (uniquely) to some $\overline{L}\in\lines$, and
\item
  distinct lines of $\goth N$ extend to distinct lines of $\goth M$.
\end{itemize}
$\goth N$ is {\em l-closed} (is a {\em  subspace} of $\goth M$) if each line of 
$\goth M$ that crosses $X$ in at least two points is an extension of a line in $\cal G$.
In that case we also say that $\goth N$ is {\em regularly} contained in $\goth M$.
Frequently, with a slight abuse of language, we refer to lines of the form $\overline{L}$
($L\in{\cal G}$) as to {\em sides} of $\goth N$.
\par\noindent
${\goth N}$ is {\em p-closed} if its sides do no intersect outside $\goth N$ i.e. 
when the following holds:
\begin{ctext}
  $L_1,L_2\in{\cal G},\; L_1\neq L_2,\; p\in\overline{L_1}\cap \overline{L_2}
  \implies p \in X$
\end{ctext}
A graph contained and p-closed in $\goth M$ is said to be {\em freely contained in} $\goth M$.
In what follows the phrases
\begin{itemize}\def\labelitemi{}\itemsep-2pt
\item {\em $X$ is a $K_n$-graph (freely) contained in $\goth M$} and
\item {\em $K_X$ is (freely) contained in $\goth M$}
\end{itemize}
will be used equivalently, together with their (admissible) stylistic variants.


\section{Complete subgraphs freely contained in \PSTS's}

Clearly, there are configurations with freely contained subgraphs.
%
\begin{prop}\label{pelny2minSTS}[{\normalfont cf. \cite{klimczak} or \cite{pascvebl}}]
  Let $n\geq 2$ be an integer. 
  $\GrasSpace(n+1,2)$ is a $\binkonf(n,+1)$-configuration which freely contains $K_n$.
\end{prop}
\begin{prop}\label{minSTS2pelny}
  Let $n\geq 2$ be an integer.
  A smallest \PSTS\ that freely contains the complete graph $K_n$ is a 
  $\binkonf(n,+1)$-configuration.
  Consequently, it is a \BSTS.
\end{prop}
\begin{proof}
  Let ${\goth M} = \struct{S,\lines}$ freely contain $K_X = \struct{X,\sub_2(X)}$,
  $|X| = n$. Then $X \subset S$ and for each edge $e\in\sub_2(X)$ there is a third point
  $\infty_e\in\overline{e}\setminus e$ of $\goth M$ on $\overline{e}$.
  Therefore,
  $|S| \geq |X| + |\sub_2(X)| = \binom{n+1}{2}$.
  Write $X^\infty = \{ e_\infty\colon e\in\sub_2(X) \}$. The rank of a point $x\in X$
  in $\goth M$ is at least $n-1$, so the rank of a point $q \in X^\infty$ is at 
  least $n-1$ as well. 
  On the other hand, there passes exactly one side of 
  $K_X$ through a point in $X^\infty$.
  Let $b_0$ be the number of lines through a point in $X^\infty$
  distinct from the sides of $K_X$. This number is minimal when the lines in question
  contain entirely points in $X^\infty$ and then through $q\in X^\infty$ there pass: 
  one  side of $K_X$ and lines contained in $X^\infty$.
  Consequently, 
  $3 b_0 \geq |\sub_2(X)|(n-2)$. This yields $b_0 \geq \binom{n}{3}$.
  Finally, 
  $|\lines| \geq |\sub_2(X)| + b_0 = \binom{n+1}{3}$.
\end{proof}
\begin{prop}\label{pelny2horyzont}
  Let ${\goth M} = \struct{S,\lines}$ be a minimal \PSTS\ which freely contains a complete
  graph $K_X = \struct{X,{\sub_2(X)}}$ and $|X| = n$. 
  Then {\em the complement of $K_X$} i.e. the structure
  $\struct{S\setminus X,\lines\setminus\{ \overline{e}\colon e\in{\sub_2(X)} \}}$
  is a
  $\binkonf(n,0)$-configuration and a subspace of $\goth M$.
  \par
  Conversely, let ${\goth N} = \struct{Z,{\cal G}}$ be a
  $\binkonf(n,0)$-configuration
  regularly contained in $\goth M$.
  Then $S\setminus Z$ yields in $\goth M$ a complete $K_n$-graph
  freely contained in $\goth M$, 
  whose complement is $\goth N$.
\end{prop}
\begin{proof}
  By \ref{minSTS2pelny}, $\goth M$ is a $\binkonf(n,+1)$-configuration.
  The first statement was proved, in fact, in the proof of \ref{minSTS2pelny}.
  \par
  Let $\goth N$ be a suitable subconfiguration of $\goth M$. Set $X = S\setminus Z$.
  Then $|X| = n$. Through each point $z\in Z$ there passes exactly one line 
  $L_z \in \lines\setminus{\cal G}$. If $z_1\neq z_2$ then $L_{z_1}\neq L_{z_2}$,
  as otherwise $L_{z_1}\in{\cal G}$.
  We have 
  $|\lines\setminus{\cal G}| = \binom{n+1}{3} - \binom{n}{3}
  = \binom{n}{2} = |\{L_z \colon z\in Z \}|$ 
  and therefore 
  $\lines\setminus{\cal G} = \{ L_z\colon z \in Z \}$.
  Each line in $\lines\setminus{\cal G}$ contains exactly two elements of $X$;
  comparing the parameters we get that 
  $\struct{X,\lines\setminus{\cal G}}$ is a complete graph.
\end{proof}

As we learn from the proofs of \ref{minSTS2pelny} and \ref{pelny2horyzont}
each minimal partial {\sf STS} that freely contains a complete graph $K_n$ is associated 
with a labelling (the map $e\mapsto\infty_e$) of the points of a 
$\binkonf(n,0)$-configuration $\goth H$ by the elements of $\sub_2(X)$, $|X| = n$.
\par
Indeed, let ${\goth H} = \struct{Z,{\cal G}}$, $\mu\colon \sub_2(X)\longrightarrow Z$
be a bijection, and $n = |X|$. Then the configuration
$$
  K_X +_\mu {\goth H} := \struct{{X \cup Z},%
  {{\cal G} \cup \left\{ \{ a,b,\mu(\{ a,b \}) \} \colon \{a,b\}\in\sub_2(X) \right\}}}
$$
%
is a $\binkonf(n,+1)$-configuration freely containing $K_X$.

In what follows any bijection $\mu$ of the points of a configuration onto an arbitrary
set will be frequently named a {\em labelling}. 
\footnote{%
In \cite{klimczak} the term `improper points' was used instead of `labelling'. 
Now we think that this older term should not be used, as it suggests, incorrectly,
some connections with a `parallelism', something to do with `directions'.}

The above apparatus was fruitfully applied in \cite{klik:VC} to the case $n=4$
(labelling of the Veblen configuration) to classify $10_3$-configurations which freely 
contain $K_4$. 
Clearly, this method can be applied to arbitrary $n$,
though even in the next step: $n=5$ a classification ``by hand" of all the 
labellings of arbitrary $\konftyp(\binom{5}{2},3,\binom{5}{3},3)$-configuration 
by the elements of 
$\sub_2(X)$ with $|X|=5$ seems seriously much more complex, if executable.
In what follows we shall propose a way that may be applied to arbitrary $n$
and which seems (at least a bit) less involved.

\begin{cor}\label{cor:pelny2horyzont}
  Let $\goth M$ be a $\binkonf(n,+1)$-configuration.
  $\goth M$ freely contains a complete graph $K_n$ iff $\goth M$
  regularly contains a $\binkonf(n,0)$-subconfiguration.
\end{cor}

\ifnoweb
\subsection{Intersection properties}
\fi

\begin{prop}\label{prop:cross-compl2}
  Any two distinct complete $K_n$-graphs freely contained in a \linebreak 
  $\binkonf(n,+1)$-configuration
  share exactly one vertex.
\end{prop}
\begin{proof}
  Let $G_1 = \struct{X_1,\sub_2(X_1)}$, $G_2 = \struct{X_2,\sub_2(X_2)}$ be two $K_n$
  graphs freely contained in a $\binkonf(n,+1)$-configuration ${\goth M}$.

  First, we note that $X_1\cap X_2\neq\emptyset$.
  Indeed, suppose that $X_1\cap X_2 = \emptyset$. 
  Then $X_2$ is freely contained in the complement of $X_1$.
  From \ref{pelny2horyzont}, this complement is a $\binkonf(n,0)$-configuration,
  which contradicts \ref{minSTS2pelny}.

  Let $a\in X_1\cap X_2$. Since 
  the degree of $a$ in $K_{X_1}$, in $\goth M$, and in $K_{X_2}$ is $n-1$, 
  $G_1$ and $G_2$ have common sides through $a$.
  Assume that $b \in X_1\cap X_2$ for some $b \neq a$; as previously the sides of $G_1$
  and of $G_2$ through $b$ coincide. So, consider arbitrary $x\in X_1\setminus\{ a,b\}$.
  $G_1$ and $G_2$ both contain the sides $\overline{\{ a,x \}}$ and $\overline{\{ x,b \}}$,
  so $x \in X_2$. Finally, we arrive to $X_1 = X_2$.
\end{proof}
For a geometer the situation considered in \ref{prop:cross-compl2} has clear geometrical
meaning: if $a$ is the common vertex of two complete $K_n$ graphs 
$\struct{X_1,{\cal E}_1}$, $\struct{X_2,{\cal E}_2}$ 
freely contained in a $\binkonf(n,+1)$-configuration $\goth M$ then $a$ is the 
{\em perspective center} of 
two $K_{n-1}$-simplices 
$A_1 = X_1\setminus\{a\}$ and $A_2 = X_2\setminus\{a\}$.
This means: there is a bijective correspondence $\sigma_a$ between the vertices of
the simplices such that for every vertex $x$ of the first simplex the corresponding
vertex $\sigma_a(x)$ of the latter simplex lies on the line $\overline{a,x}$
through $a$ and $x$.
As we shall see, in this case also an analogue of a {\em perspective axis} can be found.
That is, there is a subspace $Z$ of $\goth M$ and a bijective correspondence $\zeta$
between the sides of the simplices $A_1$ and $A_2$
such that for each side $L$ of the first simplex the corresponding side
$\zeta(L)$ of the latter simplex crosses $L$ in a point on $Z$.
\begin{prop}\label{prop:cross-compl3}
  Let $G_i= \struct{X_i,\sub_2(X_i)}$, $i=1,2$ be two complete $K_n$-graphs
  freely contained in a $\binkonf(n,+1)$-configuration ${\goth M}=\struct{S,\lines}$,
  let $p\in X_1\cap X_2$, and ${\goth N}_i$ with the pointset $Z_i = S\setminus X_{3-i}$
  be the complement of $G_{3-i}$ in $\goth M$ for $i=1,2$ (cf. \ref{pelny2horyzont}).
  \begin{sentences}\itemsep-2pt
  \item\label{cros-compl3:1}
    ${\goth N}_i$ freely contains a complete $K_{n-1}$-graph, for each $i=1,2$.
  \item\label{cros-compl3:2}
    Each side of $G_i$ missing $p$ crosses exactly one side of $G_{3-i}$ and
    the latter misses $p$ as well.
    The intersection points of the corresponding sides form the set $Z_1\cap Z_2$.
  \end{sentences}
\end{prop}
\begin{proof}
  It seen that 
  $X_{i}\setminus\{ p \}$ is a $K_{n-1}$-graph contained in $Z_{i}$ 
  and, clearly, it is p-closed. This proves \eqref{cros-compl3:1}.
  \par
  Let $p\notin e\in\sub_2(X_i)$, 
  then $\overline{e}\setminus e$ is a point $u_e$ in $S$.
  Since $u_e\notin X_i$, we have $u_e\in Z_{3-i}$.
  Moreover, $u_e\notin X_{3-i}$, since 
  $X_{3-i} \subset \bigcup\{ \overline{p,x}\colon x\in X_i \}$,
  so $u_e\in Z_i$.
  Therefore, there is a side $e'$
  of $G_{3-i}$ (a line of ${\goth N}_{3-i}$)
  which passes through $u_e$. 
  Statement \eqref{cros-compl3:2} is now evident.
\end{proof}
Even in the smallest possible case ($n = 4$) we have a 
$\binkonf(n,1)$-configuration (the fez configuration, cf. \cite{klik:VC}) 
which contains a pair of
perspective triangles with the perspective center $p$ such that the correspondence of 
the form $\overline{ \{ a,b\} } \longmapsto \overline{ \{ \sigma_p(a),\sigma_p(b) \} }$
does not yield any perspective axis, but the triangles in question do have a perspective axis.

\ifnoweb\relax\else
\subsection{\BSTS's with a few complete graphs}

Let us say a few comments on binomial configurations which contain `a few' complete graphs.
First, as an immediate consequence of 
\ref{prop:cross-compl2} and \ref{prop:cross-compl3}\eqref{cros-compl3:1} we can construct
the following
\begin{exm}\label{exm:oneK2noK}
  {\em
  Let ${\goth N}$ be a $\binkonf(n,0)$-configuration 
  which does not freely contain any $K_{n-1}$-graph. Let $|X|=n$ and 
  $\mu\colon \sub_2(X) \longrightarrow \{ \text{the points of }{\goth N} \}$ 
  be a bijection.
  Then $K_X +_\mu {\goth N}$ is a $\binkonf(n,+1)$-configuration which freely contains
  exactly one $K_n$ graph.}
  \par\noindent
  {\sc Indeed}, if $Y$ was another $K_n$-subgraph of $K_X +_\mu {\goth N}$ 
  then $Y$ would yield a $K_{n-1}$ graph in the 
  complement of $K_X$ i.e. in $\goth N$.
  Note that there do exist $\konftyp(10,3,10,3) = \binkonf(5,0)$-configurations without
  $K_4$-subgraphs (cf. \cite{betten}).
\myend
\end{exm}
There are lots of quite tricky contructions which produce a binomial configuration
with prescribed small (=0,1,2,3) number of its frreely contained maximal complete
subgraphs, but we pass over them. 
Simply because the uniform inductive procedure
which produces such a configuration with arbitrary admissible number of complete
subgraphs exists; the procedure will be shown in \ref{thm:allKn}. 
However, these constructions, mostly involving suitable defined labelling,
enable us to say some more about the structure of the configuration we
produce, e.g. about configurrations complementary to complete subgraphs.
Let us quote one of such contructions.
%

Note that if a $\binkonf(n,+1)$-configuration $\goth M$ freely contains two $K_n$-graphs
then the two arising complementary $\binkonf(n,0)$-configurations need not be isomorphic.
\begin{exm}\label{exm:dwaKn}
  Let ${\goth N}= \struct{Z,{\cal G}}$ be a $\binkonf(n,-1)$-configuration, so 
  $|Z|=\binom{n-1}{2}$. Let $X$ be an arbitrary $(n-1)$-element set, and let 
  $\mu_1,\mu_2\colon \sub_2(X)\longrightarrow Z$ be two labellings.
  Finally, let $p\notin Z \cup (X \times \{ 1,2 \})$,
  $S = X\times\{ 1,2\} \cup Z \cup \{p\}$. Consider the following system $\lines$ of blocks:
  \begin{eqnarray*}
    \lines & = & {\cal G} \\
    \strut & \cup & \left\{ \{ p,(x,1),(x,2) \}\colon x \in X \right\} \\
    \strut & \cup & \left\{ \{ (x,1),(y,1),\mu_1(\{ x,y \})) \} \colon \{x,y\}\in\sub_2(X) \right\} \\
    \strut & \cup & \left\{ \{ (x,2),(y,2),\mu_2(\{ x,y \})) \} \colon \{x,y\}\in\sub_2(X) \right\}.
  \end{eqnarray*}
  Then 
  ${\goth M} := \struct{S,\lines}$ is a $\binkonf(n,+1)$-configuration. $\goth M$ freely 
  contains two $K_n$-graphs 
  $K_{Z_1}$, $K_{Z_2}$, $Z_i = \{p\} \cup (X\times\{i\})$, $i=1,2$.
  On the other hand, the subconfiguration of $\goth M$ complementary to $K_{Z_i}$ is,
  evidently, isomorphic to $K_X +_{\mu_{3-i}} {\goth N}$.

  \par
  If $n\geq 5$  
  then 
  there are labellings $\mu_1,\mu_2$ of $\sub_2(X)$, $|X|=n-1$ such that 
  $K_X +_{\mu_1}{\goth N} \not\cong K_X +_{\mu_2} {\goth N}$. 
  For $n=5$ the required labellings are listed in \cite{klik:VC}.
\myend
\end{exm}
\fi 

\begin{prop}\label{prop:cross-compl-line}
  Let $\struct{X_i,\sub_2(X_i)}$, $i=1,2,3$ be three distinct $K_n$ graphs 
  freely contained in a $\binkonf(n,+1)$-configuration $\goth M$. Let 
  $c_k\in X_i\cap X_j$ for all $\{k,i,j\} = \{1,2,3\}$.
  Then $\{c_1,c_2,c_3\}$ is a line of $\goth M$.
\end{prop}
\begin{proof}
  We have $c_3 \in X_1 \cap X_2$, so $X_2\setminus \{c_3\}$ consists of all the 
  `third points' on sides of $X_1$ through $c_3$ i.e. 
  \begin{ctext}
    $X_2\setminus \{c_3\} = \{ \overline{c_3,x}\setminus\{ c_3,x \}\colon 
    x\in X_1\setminus \{ c_3\} \}$.
  \end{ctext}
  From the assumption, $c_1\in X_2$ and thus $\{ c_1,c_3,u \}$ is a line of $\goth M$
  for some $u \in X_1\setminus\{ c_3 \}$.
  With analogous reasoning we have
  \begin{ctext}
    $X_3\setminus \{c_1\} = \{ \overline{c_1,x}\setminus\{ c_1,x \}\colon 
    x\in X_2\setminus \{ c_1\} \}$,
  \end{ctext}
  so $u\in X_3$. Finally, with \ref{prop:cross-compl2} we have $u = c_2$: the claim.
\end{proof}
\begin{cor}\label{cor:cross-compl-sides}
  Let ${\goth M} = \struct{S,\lines}$ be a $\binkonf(n,+1)$-configuration.
  Let \linebreak
  $X_1,X_2,X_3\in\sub_n(S)$ be pairwise distinct, ${\cal E}_i = \sub_2(X_i)$ and 
  $G_i = \struct{X_i,{\cal E}_i}$ for $i=1,2,3$.
  \begin{sentences}\itemsep-2pt
  \item
    Assume that $G_1$ and $G_2$ are freely contained in $\goth M$.
    Then $G_1,G_2$ have exactly $n-1$ common sides i.e.
    \begin{ctext}
      $|\left\{ \overline{e}\colon e\in{\cal E}_1 \right\} \cap 
      \left\{ \overline{e}\colon e\in{\cal E}_2 \right\}| = n-1$.
    \end{ctext}
  \item
    Assume that $G_1,G_2,G_3$ are freely contained in $\goth M$.
    Then there is exactly one side common to $G_1$, $G_2$, and $G_3$.
  \end{sentences}
\end{cor}
%

\subsection{The structure of complete subgraphs}

Next, we pass to an analysis of possible `many' subgraphs freely contained in 
a binomial configuration.
\begin{prop}\label{prop:maxK}
  Let ${G}_i = \struct{X_i,\sub_2(X_i)}$, $i=1,\ldots,m$ be a family of 
  $m$ distinct $K_n$-graphs
  freely contained in a $\binkonf(n,+1)$-configuration ${\goth M} = \struct{S,\lines}$.
  \begin{sentences}\itemsep-2pt
  \item\label{prop:maxK:1}
    Set $I = \{ 1,\ldots,m \}$.
    The map $q\colon \sub_2(I)\longrightarrow S$ determined (cf. \ref{prop:cross-compl2}) 
    by the condition
    \begin{ctext}
      $q^{i,j} = q(\{i,j\}) \in X_i \cap X_j$ for each $\{i,j \}\in\sub_2(I)$
    \end{ctext}
    embeds $\GrasSpace(I,2)$ into $\goth M$.
  \item\label{prop:maxK:2}
    Consequently, $m \leq n+1$.
  \item\label{prop:maxK:3}
    If $m=n$ then $\goth M$ freely contains one more, $(n+1)$-st $K_n$-graph. 
  \end{sentences}
\end{prop}
\begin{proof}
  Ad \eqref{prop:maxK:1}:\quad 
  From \ref{prop:cross-compl-line}, $X_i\cap X_j \cap X_k = \emptyset$ 
  for distinct $i,j,k$ in $I$ and thus the map $q$ is an injection.
  Moreover, \ref{prop:cross-compl-line} also yields that $q$ maps each line of 
  $\GrasSpace(I,2)$ onto a line of $\goth M$.
\par\noindent
  \eqref{prop:maxK:2} is immediate now.
\par\noindent
  Ad \eqref{prop:maxK:3}:\quad
  For each $i\in I$ there is exactly one point $d_i\in X_i\setminus\bigcup_{j\neq i}X_j$.
  Write $X_0 = \{ d_i\colon i\in I\}$, clearly, $|X_0| = n$.
  Let $i,j\in I$ be distinct; from definition, $q^{i,j}\neq d_i,d_j$.
  From  \ref{prop:cross-compl-line} 
  we get that
  for every $k\in I$, $k\neq i,j$ the side $\overline{q^{i,j},q^{i,k}}$ of $G_i$ crosses
  $X_j$ in the point $q^{j,k}$. So, the side $\overline{q^{i,j},d_i}$ of $G_i$ 
  crosses $X_j$ in a point distinct from all the $q^{j,k}$ i.e. in the point $d_j$.
  Thus $X_0$ is a complete graph with the sides
  $\{ d_i,d_j,q^{i,j} \}$, $\{ i,j \}\in\sub_2(I)$.
  It is seen that $X_0$ is freely contained in $\goth M$.
\end{proof}
\begin{cor}\label{cor:maxKgras}
  A $\binkonf(n,+1)$-configuration $\goth M$ freely contains $n+1$ $K_n$-graphs
  iff ${\goth M}\cong\GrasSpace(n+1,2)$.
\end{cor}
\begin{proof}
  In view of \ref{prop:maxK}\eqref{prop:maxK:1} it suffices to note that the sets 
  $S(i) = \{e\in\sub_2(I)\colon i\in e \}$ 
  are the maximal cliques of $\GrasSpace(I,2)$
  which are not lines of $\GrasSpace(I,2)$ (cf. \cite{perspect}). It is seen that
  each of them is freely contained in $\GrasSpace(I,2)$ for arbitrary set $I$
  with $|I|\geq 3$.
\end{proof}

The results obtained can be summarized in the following Proposition.
\begin{prop}\label{lem:maxK:struktura}
  Let ${G}_i = \struct{X_i,\sub_2(X_i)}$, $i=1,\ldots,m$ be a family of 
  $m$ distinct $K_n$-graphs
  freely contained in a $\binkonf(n,+1)$-configuration ${\goth M} = \struct{S,\lines}$.
  Set $I = \{ 1,\ldots,m \}$,
  $Z_i := X_i \setminus \bigcup_{k\in I\setminus\{ i \}} X_k$,
  $Z := S \setminus \bigcup_{i\in I}X_i$,
  ${\cal E}_i := \{ \overline{e}\colon e\in\sub_2(X_i) \}$,
  ${\cal G}_i := {\cal E}_i \setminus \bigcup_{k\in I\setminus \{ i \}} {\cal E}_k$,
  ${\cal G} := \lines\setminus \bigcup_{i\in I}{\cal E}_i$ for every $i\in I$,
  $q^{i,j}\in X_i,X_j$, $Q:=\{ q^{i,j}\colon \{ i,j\}\in \sub_2(I) \}$.
  \begin{sentences}\itemsep-2pt
  \item\label{maxK:struktura:1}
    If $L\in{\cal G}$ then $L \subset Z$.
  \item\label{maxK:struktura:2}
    Let $L\in\lines$. If $|L\cap Z|\geq 2$ then $L\in{\cal G}$.
  \item\label{maxK:struktura:3}
    $|Z_i| = n - m + 1$ for every $i\in I$.
  \item\label{maxK:struktura:4}
    Let $\{i,j\}\in\sub_2(I)$.
    Then $Z_i\cup \{q^{i,j} \}$ and $Z_j \cup \{ q^{i,j} \}$ are 
    two $K_{n-m+2}$-graphs  with the common sides through $q^{i,j}$. 
    Each of them is freely contained in $\goth M$.
  \item\label{maxK:struktura:5}
    $|Z| = \binom{n+1-m}{2}$
  \item\label{maxK:struktura:6}
    $|{\cal G}| = \binom{n+1-m}{3}$
  \item\label{maxK:struktura:7}
    Let $L\in{\cal G}_i$ for an $i\in I$. 
    Then $|L\cap X_i| = 2$ and $L\cap X_i \subset Z_i$, $|L\cap Z| =1$
  \item\label{maxK:struktura:8}
    Let $e\in\sub_2(Z_i)$ for an $i\in I$. Then $\overline{e}\in{\cal G}_i$.
  \item\label{maxK:struktura:9}
    $|{\cal G}_i| = \binom{|Z_i|}{2} = \binom{n+1-m}{2}$ for every $i\in I$.
  \item\label{maxK:struktura:10}
    Let $i\in I$. Through every point $p\in Z$ there passes exactly one $L\in{\cal G}_i$.
  \item\label{maxK:struktura:11}
    The structure $\struct{Z,{\cal G}}$ is a $\binkonf(n-m,+1)$-configuration
    regularly contained in $\goth M$.
  \end{sentences}
\end{prop}
\begin{proof}
\eqref{maxK:struktura:1}:\quad
  Suppose $a \in L\cap X_i$ for some $i\in I$ and $L\in{\cal G}$. 
  Comparing point-ranks we note that 
  all the lines of $\goth M$ through $a$ are the sides of $G_i$, so $L\in{\cal E}_i$:
  a contradiction.
\par\noindent
\eqref{maxK:struktura:2}:\quad
  Suppose $L\notin{\cal G}$, then there are $i\in I$ and an edge $e$ of $G_i$ such that 
  $L = \overline{e}$. Clearly, $|L\cap X_i|=2$, so $|L\cap Z|\leq 1$.
\par\noindent
\eqref{maxK:struktura:3}:\quad
  Evident: $Z_i = \{ x\in X_i\colon x \neq q^{i,k}, k\in I\setminus\{i\} \}$
  and $|I\setminus\{ i\}| = m-1$.
\par\noindent
\eqref{maxK:struktura:4}:\quad
  Evidently, any two points in $Z_i$ and any two points in $Z_j$ are on a line of $\goth M$:
  a suitable side of $G_i$ or of $G_j$ resp.
  Let $a\in Z_i$. Then 
  $\overline{q^{i,j},a}\setminus\{ q^{i,j},a \}$ is a point $b$ of $X_j$. 
  Suppose $b = q^{j,k}$  for some $k\in I$. 
  From \ref{prop:cross-compl-line}, $a=q^{i,k}$: a contradiction;
  thus $b\in Z_j$.
\par\noindent
\eqref{maxK:struktura:5}:\quad
  By \ref{prop:cross-compl2} and \ref{prop:cross-compl-line},  
  $|\bigcup_{i\in I}X_i| = m\cdot n - \binom{m}{2}\cdot 1 =: \gamma(n)$;
  we compute 
  $\binom{n+1}{2} - (\binom{m+1-n}{2} + \gamma(n))$ = 0.
\par\noindent
\eqref{maxK:struktura:6}:\quad
  Analogously, by \ref{prop:cross-compl-line} and \ref{cor:cross-compl-sides}, 
  $|\bigcup_{i\in I}{\cal E}_i| = 
  m\cdot\binom{n}{2} - \binom{m}{2}\cdot(n-1) + \binom{m}{3}\cdot 1 =:\delta(n)$,
  and then
  $\binom{n+1}{3} - (\binom{n+1-m}{3} + \delta(n)) = 0$.
\par\noindent
\eqref{maxK:struktura:7}:\quad
  It is clear that any $L\in {\cal G}_i\subset{\cal E}_i$ 
  crosses $X_i$ in a pair $a,b$ of points.
  Suppose $a\notin Z_i$. Then $a = q^{i,k}$ for some $k\in I$, $k\neq i$ and then 
  $L\in{\cal E}_i,{\cal E}_k$. This yields $a,b\in X_i$.
  Suppose $L = \{ a,b,c \}$ and $c \in X_k$ for $k\in I$. Then $L\in{\cal E}_k$, $k\neq i$,
  so $L\notin{\cal G}_i$.
\par\noindent
\eqref{maxK:struktura:8}:\quad
  Suppose $\overline{e}\notin{\cal G}_i$, so there is $k\neq i$ such that 
  $\overline{e}\in{\cal E}_k$. This means: $G_k$ contains an edge $e'$ with
  $\overline{e} = \overline{e'}$. 
  Then $e\cap e'\neq \emptyset$, so $e\cap X_k\neq \emptyset$,
  which contradicts $e\subset Z_i$.
\par\noindent
\eqref{maxK:struktura:9}:\quad
  Immediately follows form \eqref{maxK:struktura:7} and \eqref{maxK:struktura:8}.
\par\noindent
\eqref{maxK:struktura:10}:\quad
  In view of \eqref{maxK:struktura:7},
  the map ${\cal G}_i \ni L \longmapsto p\in L\cap Z$ is well defined.
  Clearly, it is injective. 
  From \eqref{maxK:struktura:9} and \eqref{maxK:struktura:3} it is also surjective, 
  and this is exactly the claim.
\par\noindent
\eqref{maxK:struktura:11}:\quad
  Immediate, after \eqref{maxK:struktura:1}, \eqref{maxK:struktura:2},
  \eqref{maxK:struktura:5}, and \eqref{maxK:struktura:6}.
\end{proof}

Let $I = \{ 1,\ldots,m \}$ be arbitrary, let $n > m$ be an integer, and let
$X$ be a set with $n-m+1$ elements.
Let us fix an arbitrary $\binkonf(n-m,+1)$-configuration 
${\goth B} = \struct{Z,{\cal G}}$.
Assume that we have two maps $\mu,\xi$ defined:
$\mu\colon I\longrightarrow Z^{\sub_2(X)}$ 
and
$\xi\colon I\times I\longrightarrow S_X$, such that
$\xi_{i,i} = \id$, $\xi_{i,j} = \xi_{j,i}^{-1}$, and
$\mu_i$ is a bijection for all $i,j\in I$.
Let $S = Z \cup (X\times I) \cup \sub_2(I)$ (to avoid silly errors we assume that
the given three sets are pairwise disjoint).
On $S$ we define the following family $\lines$ of blocks
\begin{eqnarray*}
  \lines & = & {\cal G} \\
  \strut & \cup & \text{ the lines of } \GrasSpace(I,2) \\
  \strut & \cup & \left\{ \{ \{i,j\}, (x,i), (\xi_{i,j}(x),j)  \}
                  \colon \{i,j\}\in\sub_2(I), x \in X  \right\} \\
  \strut & \cup & \left\{ \{ (a,i), (b,i), \mu_i(\{ a,b \}) \} 
                  \colon \{ a,b \}\in \sub_2(X), i\in I  \right\}.
\end{eqnarray*}
Write 
\begin{equation}\label{def:SPS}
 m \bowtie^\mu_\xi {\goth B} = \struct{S,\lines}.
\end{equation}
It needs only a straightforward (though quite tidy) verification to prove that

\begin{center}
{\em ${\goth M}:= m \bowtie^\mu_\xi {\goth B}$ is a $\binkonf(n,+1)$-configuration}. 
\end{center}

For each $i\in I$ we set $Z_i = X\times\{ i \}$, $S_i = \{ e\in\sub_2(I)\colon i\in e \}$,
and $X_i = Z_i \cup S_i$. It is seen that 
{\em $\goth M$ freely contains $m$ $K_n$-graphs $X_1,\ldots,X_m$}.
Indeed, let us write $a\oplus b = c$ when $\{a,b,c \}$ is a line (of the configuration in
question). Then we have $\{i,j\}\oplus\{i,k\} = \{j,k\}$,
$(a,i)\oplus(b,i) = \mu_i(\{a,b\})$, and $(a,i)\oplus\{i,j\} = (\xi_{i,j}(a),j)$.
It is seen that the point $\{i,j\}$ is the perspective center of two subgraphs $Z_i,Z_j$
of $\goth M$. So, we call the 
configuration $m \bowtie ^\mu_{\xi} {\goth B}$
{\em a system of perspective $(n-m+1)$-simplices}.
Define $\mu^o_i\colon\sub_2(Z_i)\longrightarrow Z$ by the formula
$\mu^o_i(\{ (x,i),(y,i) \}) = \mu(\{ x,y \})$; 
it is seen that the configuration $\goth B$ is the common `axis' of  the configurations 
$\struct{Z_i,\sub_2(Z_i)} +_{\mu^o_i} {\goth B}$
contained in $\goth M$.
\par
Note that the words `perspective', `axis', and `simplices' are used to suggest some 
{\em formal} similarities to objects considered in geometry.
Such a usage does not mean that the considered binomial configuration 
$m \bowtie^mu_\xi {\goth B}$ is necessarily realizable in a desarguesian projective
space.

Let us consider two special cases
of the above definition of a system of perspective simplices.
\begin{sentences}\itemsep-2pt
\item
  Let $m = n-1$. Then $\goth B$ consists of a single point $p$: the {\em center} of $\goth M$.
  Consequently, $\mu_i$ is constant, $\mu_i \equiv p$. Moreover, the set $X$ which 
  appears in the definition has 2 elements and then $|S_X| =2$. Then instead of a map $\xi$
  one can consider a graph ${\cal P}\subset\sub_2(I)$ defined by
  $\{ i,j\} \in {\cal P} \iff \xi_{i,j} = \id$. And then 
  {\em the system $m \bowtie^p_\xi \struct{\{ p \},\emptyset}$
  of $m$ perspective segments
  (of 2-simplices) is the multiveblen configuration 
   $\xwlepp({},{m},{\cal P},{},{\GrasSpace(m,2)})$}  
  (cf. \cite{pascvebl}, \cite{mveb2proj}).
\item
  Let $m = n-2$. Then $\goth B$ consists of a single $3$-point line, $L = \{a,b,c\}$.
  Up to a permutation of $X$ there is a unique bijection $\mu\colon\sub_2(X)\longrightarrow L$.
  Finally 
  {\em the system of $m$ perspective triangles 
  $m \bowtie^\mu_\xi \struct{L,\{ L \}}$
  is the system of triangle perspectives
    $\xwlepq({I},{},{\xi},{},{\GrasSpace(I,2)})$}
  (cf. \cite{linstp}).
\end{sentences}

Now, till the end of this section writing
`a configuration contains $m$ graphs' we mean
`a configuration contains {\em at least} $m$ graphs'.
\begin{thm}\label{thm:SPS}
  Let ${\goth M}$ be a $\binkonf(n,+1)$-configuration.
  The following conditions are equivalent.
  \begin{enumerate}[(i)]\itemsep-2pt
  \item
    $\goth M$ freely contains $m$ $K_n$-graphs.
  \item
    $\goth M$ is a system of $m$ perspective $(n-m+1)$-simplices i.e. 
    ${\goth M} \cong m \bowtie^\mu_\xi {\goth B}$ for a 
    $\binkonf(n-m,+1)$-configuration $\goth B$ and some (admissible) maps $\mu,\xi$.
  \end{enumerate}
\end{thm}
\begin{proof}
  We have already noticed that $m \bowtie^\mu_\xi {\goth B}$ contains required subgraphs.
  \par
  Let ${\goth M} = \struct{S,\lines}$ and
  let $X_1,\ldots,X_m\in\sub_n(S)$ be pairwise distinct. Assume that 
  $G_i = \struct{X_i,\sub_2(X_i)}$ is freely contained in $\goth M$ for every 
  $i = 1,\ldots,m$.
  Let us adopt the notation of \ref{lem:maxK:struktura}.
  
  Set ${\goth B} = \struct{Z,{\cal G}}$.
  Let us fix a $(n-m+1)$-element set $X$ and let 
  $\nu_i\colon Z_i \longrightarrow X$ be a fixed bijection for each $i\in I$.
  Let $a,b, \in X$ and $i,j\in I$.
  Define 
  \begin{ctext}
    if $a\neq b$ then 
    $\mu_i(\{ a,b \}) = \overline{\nu_i(a)\nu_i(b)}\setminus \{ \nu_i(a),\nu_i(b) \}$,
  \end{ctext}
  $x_{i,i} = \id_X$, and 
  \begin{ctext}
    if $i\neq j$ then
    $\xi_{i,j}(a) = b$ iff $\{q^{i,j},\nu_i(a),\nu_j(b)\}\in\lines$.
  \end{ctext}
  Finally, we define on the points of $\goth M$ the following map $F$:
  $$
    F \colon \left\{
    \begin{array}{rcl}
      Q\ni q^{i,j} & \longmapsto & \{i,j\} \\
      Z_i \ni x & \longmapsto & (x,i)  \\
      Z \ni a & \longmapsto & a
    \end{array} 
    \right.
  $$
  It is a standard student's exercise to compute that $F$ is an isomorphism of $\goth M$ and 
  $m \bowtie^\mu_\xi {\goth B}$.
\end{proof}
\begin{cor}\label{cor:maxK:pod}
  Let $\goth M$ be a $\binkonf(n,+1)$-configuration.
  \begin{enumerate}[(i)]\itemsep-2pt
  \item
    $\goth M$ freely contains $n-1$ graphs $K_n$ iff $\goth M$ is (isomorphic to)
    a multiveblen configuration.
  \item
    $\goth M$ freely contains $n-2$ graphs $K_n$ iff $\goth M$ is (isomorphic to)
    a system of triangle perspectives.
  \end{enumerate}
\end{cor}

Particular instances of \ref{cor:maxKgras} and \ref{cor:maxK:pod} in case $n=4$ can be 
found in \cite{klik:VC}: 
a $10_3$ configuration contains four $K_4$ iff it contains five $K_4$
iff it is a Desargues Configuration; 
a $10_3$ configuration contains three $K_4$ iff it is a 
multiveblen configuration i.e. iff 
it is the Desargues or it is the Kantor $10_3 G$-configuration (cf. \cite{kantor});
a $10_3$ configuration contains two $K_4$ iff it is a system of triangle perspectives
i.e. 
it is one of the following: the Desargues, the Kantor $10_3G$, or the fez configuration.

\section{Existence problems}

\begin{prop}\label{jedenup}
  If there is a $\binkonf(n,0)$-configuration which freely contains exactly
  $m$ maximal complete subgraphs where $m\leq n-2$ then there exists
  a $\binkonf(n,1)$-configuration which freely contains exactly $m+1$ 
  maximal complete subgraphs.
\end{prop}
\begin{proof}
  Let $Y_1,\ldots,Y_m$ be the $K_{n-1}$-subgraphs of a 
  $\binkonf(n,0)$-configuration ${\goth M} = \struct{S,\lines}$.
  Set $I = \{ 1,\ldots,m \}$.
  Let us reprezent $\goth M$ as a system of perspectives, so
  let $q_{i,j} \in Y_i \cap Y_j$ for distinct $i,j$ and 
  $Q= \{ q_{i,j}\colon \{i,j\}\in \sub_2(I) \}$, 
  $G_i = Y_i\setminus Q$, and let $Z$ be an ``axis" i.e. the intersection
  of all the complementary subconfigurations of the $Y_i$'es.
  Let $X$ be an arbitrary set disjoint with $S$ of cardinality $n$
  and let $P\in\sub_m(X)$. Let us number the elements of $P$:
  $P = \{ p_1,\ldots,p_m \}$ and the elements of $X\setminus P$:
  $X\setminus P = \{ x_1,\ldots,x_{n-m} \}$. Note that $n-m\geq 2$.
  For each $\{i,j\} \in \sub_2(I)$ we introduce the triple 
  $\{ p_i,p_j,q_{i,j} \}$ as a new line. 
  The number of points in each of the sets $G_i$ is $n-m$; let 
  $\sigma_i$ be an arbitrary bijection of $G_i$ onto $X\setminus P$.
  The second family of new lines consists of the triples
  $\{ p_i,x,\sigma_i(x) \}$ with $x\in G_i$, $i\in I$.
  Finally, the third class of the new lines consists of the triples
  $\{ x_i,x_j,\mu(x_i,x_j) \}$, where $\mu$ is an arbtrary labelling of
  the edges of the graph $K_{X\setminus P}$ by the elements of $Z$:
  it is possible due to cardinalities of the sets in question.
  Let ${\goth M}^*$ be the structure defined on the point universe
  $S\cup X$, whose lines are the lines of $\goth M$ and the three classes
  of new lines introduced above. It is seen that ${\goth M}^*$ is
  a $\binkonf(n,1)$-configuration. It is also evident that $K_X$ and
  $K_{Y_i\cup\{ p_i \}}$ for $i\in I$  are $K_n$-subgraphs freely contained in
  ${\goth M}^*$.
  Suppose that ${\goth M}^*$ contains another freely contained $K_n$-subgraph $K_Y$,
  let $p\in X\cap Y$. Then $Y_0:= Y\setminus X = Y\setminus \{ p \}$ is a 
  $K_{n-1}$ subgraph of $\goth M$. Consequently, $Y_0 = Y_i$ for some $i\in I$.
  Suppose that $p\neq p_i$, then the two subgraphs $K_{Y_i\cup\{ p_i \}}$
  and $K_Y$ freely contained in ${\goth M}^*$ have more than a point in common.
  Consequently, $p=p_i$ and $Y = Y_i\cup\{ p_i \}$.
  Thus ${\goth M}^*$ freely contains exactly $m+1$ $K_n$-subgraphs.
\end{proof}

\begin{prop}\label{dwadown}
  If there exists a $\binkonf(n,+1)$-configuration which freely contains exactly two
  $K_n$ graphs then there is also a $\binkonf(n,1)$-configuration
  without any $K_n$-subgraph freely contained in it.
\end{prop}
\begin{proof}
  \def\Mx{{\goth M}^\ast}
  Let $n \geq 4$. Let ${\goth M} = \struct{S,\lines}$ be a $\binkonf(n,+1)$-configuration
  which freely contains exactly two complete $K_n$-graphs $X_1,X_2$.
  Let $p\in X_1\cap X_2$. Set $A_i = X_i\setminus\{ p \}$.
  Let $e_1\in\sub_2(A_1)$ and $e_2\in\sub_2(A_2)$ such that 
  $\overline{e_1}\cap\overline{e_2}\ni q$ for a point $q$ 
  (cf. \ref{prop:cross-compl3}\eqref{cros-compl3:2})
  Let $e_1 = \{a_1,b_1\}$, $e_2 = \{a_2,b_2\}$ such that 
  $b_2\notin\overline{p,a_1}$ and $a_2\notin\overline{p,b_1}$.
  We replace the two lines
  $\{ q,a_1,b_1 \}$ and $\{ q,a_2,b_2 \}$ of $\goth M$
  by two other triples
  $\{ q,a_1,b_2 \}$ and $\{ q,b_1,a_2 \}$;
  let $\Mx$ be the obtained incidence structure. 
  Clearly, $\Mx$ is a $\binkonf(n,+1)$-configuration.
  {\em $\Mx$ does not freely contain any $K_n$-graph}
  \par\noindent{\sc Indeed}:
  Suppose that $\Mx$ freely contains a $K_n$-graph $Y$.
  Let us have a look at the collinearity graph $A_{\goth K}$ of an arbitrary configuration
  $\goth K$.
  Clearly, if $K_X$ is contained in $\goth K$ then $K_X$ is a subgraph of $A_{\goth K}$.
  In our case exactly two edges $e_1,e_2$ of $A_{\goth M}$ were replaced by two (other) edges
  $\{ a_1,b_2 \}$, $\{ a_2,b_1 \}$ to form $A_{\Mx}$.
  So, $K_Y$ cannot be build entirely from the edges missing $e_1\cup e_2$.
  \par
  (1) $Y\cap e_1 \neq \emptyset \neq Y\cap e_2$. Indeed, suppose, eg. $Y\cap e_1 = \emptyset$.
  The same pairs of points in $S\setminus e_1$ (except $e_2$) are collinear in 
  $\goth M$ and in $\Mx$ and therefore $Y$ is a $K_n$-graph in $\goth M$: a contradiction.
  \par
  (2) $e_1\not\subset Y$ and $e_2\not\subset Y$: the pair of points in $e_1$ is not
  collinear in $\Mx$, and, analogously the points in $e_2$ are also not collinear.
  \par
  So, $Y$ contains exactly one point $x_1$ in $e_1$ and one point $y_2$ in $e_2$.
  Clearly, $x_1,y_2$ must be collinear in $\Mx$.
  \par
  (3) Suppose that $y_2 \in \overline{p,x_1}$. Without loss of generality we can 
  take $x_1 = a_1$, $y_2 = a_2$. Then $b_1,b_2\notin Y$.
  For points in $S\setminus \{ b_1,b_2 \}$ exactly the same pairs are collinear in 
  $\goth M$ and in $\Mx$, so $Y$ is a subgraph of $\goth M$, which is impossible.
  \par
  Without loss of generality we can assume that $a_1,b_2 \in Y$ and $a_2,b_1\notin Y$.
  The following three cases should be considered:
  \begin{enumerate}[(1)]\setcounter{enumi}{3}\itemsep-2pt
  \item\label{bzz1}
    $a_2 \in \overline{p,a_1}$, and $b_2 \in \overline{p,b_1}$,
  \item\label{bzz2}
    $a_2 \in \overline{p,a_1}$ and $b_2 \notin \overline{p,b_1}$,
  \item\label{bzz3}
    $a_2 \notin \overline{p,a_1}$ and $b_2 \notin \overline{p,b_1}$.
  \end{enumerate}
  \par\eqref{bzz1}:
    Note that the sides of $Y$ and the lines of $\Mx$ through vertices of $Y$ coincide. 
    So, $\{ p, a_1\}$ is an edge of $Y$ and thus $p\in Y$.
    Take any point $c_1\in A_1\setminus e_1$.
    Then $c_1,b_2$ are not collinear in $\Mx$, so $c_1\notin Y$. 
    Let $c_2\in \overline{p,c_1}\setminus\{ p,c_1 \}$ (this line of $\goth M$ 
    was unchanged), then $c_2\in Y$.
    But $c_2$ and $a_1$ are not collinear in $\Mx$ and a  contradiction arizes.
  \par\eqref{bzz2}:
    In this case also necessarily $p\in Y$. 
    Let $c_2\in\overline{p,b_1}\setminus\{ p,b_1 \}$; then $c_2\neq b_2$ and $c_2\in Y$.
    But, contradictory, $a_1,c_2$ are not collinear in $\Mx$.
  \par\eqref{bzz3}:
    Now, either $p\in Y$ or $c_2 \in Y$,
    where $\{p,a_1,c_2\}\in\lines$ ($c_2\in X_2$). 
    If $p\in Y$ then we take 
    $c_1\in \overline{p,a_2}\setminus\{ p,a_2 \}$;
    then $c_1\in Y$. An inconsistency appears, as $c_1,b_2$ are not collinear in $\Mx$.
    Consequently, $c_2\in Y$ and $p\notin Y$.
    Therefore, the point $c_1$ in $\overline{p,b_2}\setminus\{p,b_2\}$ is in $Y$ 
    ($c_1 \in X_1$). But $c_1,c_2$ are not collinear in $\Mx$, though.
\end{proof}

As an important consequence we obtain now
\begin{thm}\label{thm:allKn}
  Let $m,n$ be integers, $4\leq n$, and $1\leq m \leq n-1$ or $m=n+1$.
  Then there exists a $\binkonf(n,+1)$-configuration which freely contains
  exactly $m$ $K_n$-graphs.
\end{thm}
\begin{proof}
 Let $n\geq 4$ be an integer and let 
 $J(n)$ be the set of integers $m$ such that there is a $\binkonf(n,1)$- configuration
 with  exactly $m$ freely contained subgraphs $K_n$.
 From \ref{prop:maxK}, $J(n)\subseteq \{ 0,1,\ldots,n-1,n+1 \} =: F(n)$.
 We need to prove that $J(n) = F(n)$ for every integer $n\geq 4$.
 Clearly, this equality holds for $n=4$ (cf. \cite{klik:VC}).
 \par
 Assume that $J(n) = F(n)$ holds for an integer $n$. 
 From \ref{jedenup} and  \ref{cor:maxKgras}, 
 $J(n+1)  \supseteq \{ 1,2,\ldots,n,n+2 \}$.
 Then from \ref{dwadown} we get $0\in J(n+1)$ and therefore
 $J(n+1) = F(n+1)$. By induction, we are done.
\end{proof}
\begin{rem}
  There is no reasonable $\binkonf({0},2)$-configuration,
  there is exactly one $\binkonf({0},3)$-configuration: a line
  $\GrasSpace(3,2)$, with exactly $3$ freely contained copies of $K_2$,
  and there is exactly one $\binkonf({0},4)$-configuration:
  the Veblen configuration $\GrasSpace(4,2)$, 
  which freely contains 4 copies of $K_3$. 
\end{rem}

\section{Other known examples}

\subsection{Combinatorial Veronesians}

Let us adopt the notation of \cite{combver}.
Let $|X| = 3$.
Then the combinatorial Veronesian $\VerSpace(X,k)$ is 
a $\binkonf(k,+2)$-configuration; 
its point set is the set $\msub_k(X)$ of the $k$-element multisets with 
elements in $X$.
The maximal cliques of $\VerSpace(X,k)$ were established in 
\cite{veradjac}. From that results we learn that
\begin{fact}
  The $K_{k+1}$ graphs freely contained in $\VerSpace(X,k)$ are the sets
  $X_{a,b} := \msub_k(\{ a,b \})$, 
  $X_{b,c} := \msub_k(\{ b,c \})$, and 
  $X_{c,a} := \msub_k(\{ c,a \})$.
  Its axis is the set $X^k$. For $u\in\sub_2(X)$ the $\binkonf(k,+1)$-subconfiguration
  of $\VerSpace(X,k)$ complementary to $X_u$ 
  (with the universe $y\msub_{k-1}(X)$, $y\in X\setminus u$) 
  is isomorphic to $\VerSpace(X,k-1)$.
\end{fact}
\begin{cor}
  A $\binkonf(k,+1)$-Veronesian with $k>2$ 
  contains exactly three complete
  $K_{k}$-graphs freely contained in it.
\end{cor}

\subsection{Quasi Grassmannians}

Let us adopt the notation of \cite{skewgras}.
The configuration ${\goth R}_n$ is a $\binkonf(n,+2)$-configuration.
Recall the role of the set $X = \{1,2 \}$ if $n$ is even and 
$X= \{ 0,1,2 \}$ if $n$ is odd. 
Namely, let us cite after \cite{skewgras} the following
\begin{fact}
  The complete $K_{n+1}$-graphs freely contained in ${\goth R}_n$ are the sets
  $S(i) = \{a\in\sub_2(Y)\colon i\in a\}$, where $\sub_2(Y)$ is the point set
  of ${\goth R}_n$, $X\subset Y$, and $i \in X$.
\end{fact}
\begin{cor}
  Let $\goth M$ be a $\binkonf(n,0)$-quasi-Grassmannian, $n>2$. 
  If $n$ is even then $\goth M$
  freely contains exactly two $K_{n-1}$-graphs, and it freely contains exactly three
  $K_{n-1}$-graphs when $n$ is odd.
\end{cor}
Besides, the above indicates one more similarity between 
combinatorial Veronesians and quasi Grassmannians represented
as a fan of configurations $10_3G$.  



\bigskip
\begin{small}
\noindent
Authors' address:\\
Ma{\l}gorzata Pra{\.z}mowska, Krzysztof Pra{\.z}mowski\\
Institute of Mathematics, University of Bia{\l}ystok\\
ul. Akademicka 2, 15-246 Bia{\l}ystok, Poland\\
{\ttfamily malgpraz@math.uwb.edu.pl},
{\ttfamily krzypraz@math.uwb.edu.pl}
\end{small}

\end{document}
